### v1

# Augmented Object Intelligence: Making the Analog World Interactable with XR-Objects

Abstract.
Seamless integration of physical objects as interactive digital entities remains a challenge for spatial computing. This paper introduces Augmented Object Intelligence  (AOI), a novel XR interaction paradigm designed to blur the lines between digital and physical by equipping real-world objects with the ability to interact as if they were digital, where every object has the potential to serve as a portal to vast digital functionalities. Our approach utilizes object segmentation and classification, combined with the power of Multimodal Large Language Models (MLLMs), to facilitate these interactions. We implement the AOI concept in the form of XR-Objects, an open-source prototype system that provides a platform for users to engage with their physical environment in rich and contextually relevant ways. This system enables analog objects to not only convey information but also to initiate digital actions, such as querying for details or executing tasks. Our contributions are threefold: (1) we define the AOI concept and detail its advantages over traditional AI assistants, (2) detail the XR-Objects  system’s open-source design and implementation, and (3) show its versatility through a variety of use cases and a user study.
mixed reality, extended reality, augmented reality, augmented objects, spatial computing, object interactions, user interfaces, context menus
copyright: nonejournalyear: conference: ; ; booktitle: ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Mixed / augmented realityccs: Computing methodologies Intelligent agents
Figure 1. XR-Objects allows users to (a) select and interact with real-world objects in AR as if they were digital objects. (b) It uses automatically generated object-based context menus in AR, and objects provide information about themselves, such as a food product. (c, d) The user asks a question about the boiling time, and (e) uses the result to interact with other objects, such as (f) setting up a spatial timer on the pot with the right cooking time. (g) The timer widget is automatically placed on the object in 3D space.
[…]
(a) A pack of pasta and a pot on a stove. (b) A menu panel overlaid on the pasta, which says: ”Filiz Burgu Pasta. (c) The interface says ”listening”. (d) ”how much time do I need to cook this pasta?” (e) Cook for 10 minutes” (f) The user taps ”timer” (g) A timer on the pot, anoter one on the teapot.
Figure 2. Example of an interaction with a conventional multimodal AI assistant. While the model clearly has the capacity to produce reasonable scene understanding when an image and a prompt is provided as input, it fails in providing an anchored output that ties back to the original multimodal prompt.
[…]
Left: The user prompt says: ”Which of these products contain lactose?” and there is a photo with several beverages. Right: The LLM gives a text answer with 3 product names. The user then asks: ”Can you mark those on the image directly? I don’t speak Japanese. The LLM responds: ”Unfortunately, I cannot - my ability to interact with the physical world is limited.
1. Introduction
Modern Extended Reality (XR) platforms come with a plethora of sensors, cameras, and advanced computer vision techniques to seamlessly blend virtual content with the physical world through color passthrough and scene understanding. However, despite these technological steps, the integration of real objects into the XR environment remains somewhat superficial, treating the physical world largely as a mere backdrop rather than an interactive component. In contrast, projects like RealityCheck (Hartmann et al., 2019) or Remixed Reality (Lindlbauer and Wilson, 2018) show a future where digital and physical worlds could be closely intertwined together. Similarly, advancements in artificial intelligence (AI) are laying the groundwork for such a future, with breakthroughs in real-time unsupervised segmentation (Tian et al., 2023) combined with in-painting (Xiang et al., 2023) or generative AI content generation (Kerbl et al., 2023).
The wide availability of machine learning (ML) and computer vision technologies has also led to features that enhance digital interaction with the physical world at our fingertips. Tools like image-based search in Google Lens [^1] Google Lens: https://lens.google and utility-focused Augmented Reality (AR) features in smartphones, such as text copy & paste and real-time translation, are becoming increasingly common. Together, these tools form building blocks that could bring us a little bit closer to a future in which a total understanding of the world and its objects can be applied to our everyday interactions in XR.
In this paper, we explore a novel interaction paradigm we term Augmented Object Intelligence (AOI) that would allow any real object identified by the XR system to reveal digital data associated with the analog object and perform context-appropriate digital actions in a meaningful way. Our system XR-Objects  embodies this idea and aims to demonstrate and investigate ”semantic equality” between real and virtual objects.
Imagine a scenario as familiar as right-clicking a digital file to open its context menu, but applied to physical objects within XR — such as right-clicking on potatoes or pasta in a pot to start a cooking timer set to the correct duration, or filtering for gluten-free products on a grocery shelf through an XR interface (see Figure 1).
The leap towards physical awareness in XR-Objects represents an advancement over traditional AR, which often relies on manual input or the use of physical tracking markers. We leverage developments in spatial understanding via technologies such as Simultaneous Localization and Mapping (SLAM), available in ARCore [^2] Google ARCore: https://developers.google.com/ar and ARKit [^3] Apple ARKit: https://developer.apple.com/augmented-reality/arkit, and machine learning models for object segmentation and classification (COCO (Lin et al., 2015) via MediaPipe (Lugaresi et al., 2019)). These technologies enable us to implement AR interactions with semantic depth. We also integrate a Multimodal Large Language Model (MLLM) into our system, which further enhances our ability to automate the recognition of objects and their specific semantic information within XR spaces.
Our contributions are threefold:
  • We introduce the concept of Augmented Object Intelligence (AOI), a paradigm shift towards seamless integration of real and virtual content in XR.
  • We detail the open-sourced design and implementation of XR-Objects, our prototypical system that exemplifies AOI, alongside an exploration of diverse use cases to demonstrate its potential.
  • We provide a comparison between standard prompt-based LLM interfaces and our AOI approach for contextual information retrieval and object-centric interaction.
We demonstrate our prototype that integrates these AR and AI components in a seamless way, which is implemented for smartphones to provide access to a broad audience, as commercial headsets do not yet give programmatic access to the user’s camera stream. By open-sourcing our system, we aim to foster further innovation in the field, ultimately bringing us closer to a future where the physical and digital realms interact seamlessly, embodying the concept of Programmable Reality (for anonymity during review process, 2024) for everyday users.
2. Related Work
This section provides an overview of previous work in the realms of user interface (UI) design principles, the interaction with real-world objects, and advancements in AR technology (Speicher et al., 2019). These areas form the foundation upon which our research builds, which aims to enhance the interaction between users and the physical world through XR-Objects.
2.1. Fundamentals of UI
The challenge of bridging the cognitive gap between human users and computational systems has been a central theme in HCI. Traditional approaches have employed various layers of representations to mediate this interaction, manifesting most recognizably in the UIs of devices ranging from PCs and smartphones to IoT devices and automotive systems (Tidwell et al., 2020). Despite these advancements, the resurgence of command-line interface (CLIs) in contemporary AI interactions, as seen in the usage of prompts with large language models (LLMs) (Suh et al., 2024), suggests a potential oversimplification of user interaction paradigms. This regression shows the necessity of reevaluating our approach to UI design in the age of AI and spatial computing (Suzuki et al., 2023), where the intricacy of human intentions and the computational interpretation thereof demand a more nuanced form of representation. Figure 2 demonstrates that the multimodal AI assistant clearly has the capacity to produce reasonable scene understanding when an image and a prompt is provided as input but it fails in providing an anchored output that ties back to the original multimodal prompt.
Significantly, the role of UIs extends beyond mere facilitation of interaction, which shapes the user’s ability to navigate, understand, and command the computational system. Effective UIs support essential cognitive functions such as memory, discovery, and articulation (Blackwell, 2006), thus becoming a key factor in the widespread adoption and utility of AI and spatial computing technologies.
In addressing these challenges, our work revisits the utility of context menus — a familiar paradigm in desktop computing (Airth, 1993; Zeng and Zhang, 2014; Banovic et al., 2011) — and explores their potential in fostering familiar interactions with physical objects within XR environments. Previous demonstrations (Lepinski et al., 2009) showed the potential of such ubiquitous context menus through simulated digital notes via projection. In this work, we show a fully functional prototype using AR.
2.2. Blended Reality
The exploration of blended-reality systems is a rapidly developing field within HCI, marked by recent research that integrates real and virtual environments (Hartmann et al., 2019; Lindlbauer and Wilson, 2018; Lindlbauer et al., 2016). While a comprehensive literature review (Auda et al., 2023) exceeds the scope of this section, we draw attention to key themes that emerged within the broader XR literature.
Manipulating Perception in XR: Gonzalez-Franco and Lanier (Gonzalez-Franco and Lanier, 2017) explored how perceptual manipulations can be effectively modeled within VR environments to enrich user experience. Bonnail et al. explore how XR can leverage human memory limitations to influence user perception and behavior (Bonnail et al., 2023). Concurrently, causality-preserving asynchronous reality enables users to interact with events in a causally accurate manner despite temporal delays (Fender and Holz, 2022). Tseng et al. highlighted the risks associated with such manipulations (Tseng et al., 2022) and proposed mitigations against their malicious use (Tseng, 2023).
Interacting with Real-World Affordances: As XR platforms become increasingly accessible, enhancing physical-aware interactions can significantly elevate user experiences. DepthLab (Du et al., 2020) exemplifies advancements in real-time 3D interactions with the real world using depth maps. InteractionAdapt (Cheng et al., 2023) optimizes VR workspaces for efficient interaction across diverse physical environments, demonstrating the flexibility and adaptability of VR systems to support varied user activities.
Balancing Immersion and External Awareness: Critical to blended-reality systems is managing the balance between immersion in the virtual world and awareness of the real world. Studies such as those by Kudo et al. emphasize the importance of smoothly transitioning users and devices between realities, while also maintaining bystander awareness (Kudo et al., 2021). Further guidelines on this balance are discussed in works by Gonzalez-Franco et al. for harmonizing user experience across dimensions (Gonzalez-Franco and Colaco, 2024; Gonzalez et al., 2024).
2.3. Detecting and Interacting with AR Objects
Despite the considerable advances in AI’s capability to generate and understand complex content, our physical world remains predominantly analog, with only a fraction of daily activities and tools being enhanced by digital technology (Rogers et al., 2020). This analog nature of human experiences, from basic needs fulfillment to complex task execution, presents a significant challenge in integrating digital intelligence in a manner that feels both natural and easy to grasp.
Figure 3. The landscape of physical object interactions in XR classified across two dimensions: anchoring and content.
[…]
A diagram with two axes: The x-axis says ”arbitrary content vs. object-focused content”. The y-axis: ”manual anchoring” vs ”seamless anchoring”. There are related work groups in this 2D space, including visible marker, invisible marker, tangible input, tangible haptic proxy, digital app, digital objects.
Several research efforts have explored ways to bridge this gap by leveraging XR technologies to enable interaction with real-world objects. Figure 3 shows the landscape of physical object interactions in XR classified across two dimensions: anchoring (manual vs. seamless) and content (arbitrary vs. object-focused).
Some approaches rely on pre-registration of objects or manual setups to achieve tangible input (Du et al., 2022; Monteiro et al., 2023; Zhu et al., 2022) or tangible haptic proxy (He et al., 2023; Jain et al., 2023; Hettiarachchi and Wigdor, 2016). In contrast to these manual processes, other works utilized markers to more automatically execute the object detection and AR content anchoring. For instance, researchers embedded visible (Rajaram and Nebeling, 2022; Henderson and Feiner, 2008; Chen et al., 2020; Li et al., 2019) or invisible markers (Dogan et al., 2022, 2023a, 2023b) to documents and 3D objects for AR purposes. However, these methods often impose limitations on the types of objects that can be interacted with, requiring specific fabrication or preparation.
Our exploration acknowledges the complexity of translating digital interactions to physical objects and aims to bridge this divide by enhancing any physical object with digital functionality and contextual interaction capabilities.
Once the objects have been identified and localized by the AR platform, an important dimension is what content will be shown at the identified locations (Caetano and Sra, 2022). Previous work has focused on optimizing the presentation of digital content (e.g., existing app windows or widgets) on or around objects (Cheng et al., 2021; Han et al., 2023; Lindlbauer et al., 2019), while others have investigated the use of tangible interaction with physical objects as a means to control such digital content (Zhu et al., 2022; Monteiro et al., 2023; Suzuki et al., 2020). EditAR (Chidambaram et al., 2022) suggested capturing users and their interactions with objects nearby to create digital twins for later consumption in XR. ProcessAR (Chidambaram et al., 2021) captures instruction demonstrations related to domain-specific objects by experts, so they can be viewed by novices in situ.
InfoLED (Yang and Landay, 2019), LightAnchors (Ahuja et al., 2019), and Reality Editor (Heun et al., 2013) took this one step further to show truly object-specific content in AR, however, this was specifically for electronic objects, such as their showing their battery level or device status in AR. Researchers further suggested augmenting visualizations with dynamic AR content (Chen et al., 2020; Chulpongsatorn et al., 2023a, b; Xiao et al., 2022), however, such dynamic content was limited to printed documents or displays.
Our work builds upon these prior efforts by proposing a system that enables object-specific content and interactions for a much wider range of objects, regardless of their physical capabilities or pre-configured markers. We leverage advancements in spatial understanding via techniques like SLAM (Davison, 2003), available in ARCore and ARKit, and machine learning models for object segmentation and classification to achieve this. This allows us to implement AR interactions with semantic depth, enabling contextually relevant information and actions for any object in the user’s environment.
Figure 4. The XR-Objects  processing pipeline combines MediaPipe and ARCore for object detection and spatial tracking, respectively, integrates an MLLM for object-specific metadata retrieval and interaction, and renders UI content in 3D space.
[…]
The RGB image is processed for both MediaPipe labeling and ARCore depth map generation. The object metadata is retrieved from the MLLM. A raycast is used for 3D localization of the object. Finally, the menu content is anchored in 3D.
3. XR-Objects Implementation
XR-Objects  leverages developments in spatial understanding via tools such as SLAM, available in ARCore [^4] Google ARCore: https://developers.google.com/ar and ARKit [^5] Apple ARKit: https://developer.apple.com/augmented-reality/arkit, and machine learning models for object segmentation and classification (COCO (Lin et al., 2015) via MediaPipe (Lugaresi et al., 2019)), which enables us to implement AR interactions with semantic depth. We also integrate a Multimodal Large Language Model (MLLM) into our system, which further enhances our ability to automate the recognition of objects and their specific semantic information within XR spaces.
Platform
Given the current technological constraints of AR headsets, particularly their limited developer access to real-time camera streams, we consciously targeted smartphones for our mobile prototype development. We aim to enable anyone to try out our open-source project on their phone. Modern smartphones use similar types of ARM-based mobile chipsets as AR headsets, yielding comparable performance for real-time computer vision tasks, while providing unrestricted access to their high-resolution cameras. This enables our application to identify objects in the user’s environment and overlay digital information directly onto the physical world through the phone’s display thanks to ARCore and ARKit.
Multimodal Interaction
At the heart of XR-Objects  is a multimodal large language model (MLLM) (Yin et al., 2024) as well as a speech recognizer, which facilitate a rich interaction layer between the user and the objects. This model not only recognizes objects but also fetches and provides contextual information and actions relevant to the selected object. By integrating voice and visual inputs, our system offers a seamless and familiar interface for users to engage with their surroundings in novel ways.
3.1. Design Considerations
In developing the AOI paradigm for XR environments, we considered a range of design choices to enhance user interaction and system performance. These considerations are grounded in our review of related work and guided by our goal to seamlessly integrate digital functionalities with physical objects. Here, we explain our rationale behind key design decisions, contrasting them with alternative approaches and situating them within the broader discourse of HCI and AR.
3.1.1. Object-Centric vs. App-Centric Interaction
Traditional AR interactions often follow an app-centric model, where users must first open a specific application to access digital functionalities. In this model, users have to navigate through the app’s interface to select categories or objects of interest, and in some cases, even upload pictures for analysis. Examples of app-centric interactions include standard ChatGPT-style interfaces, where users input a query and an image, and Google Lens, which requires users to open the app and manually select the objects they wish to interact with.
In contrast, our system prioritizes an object-centric approach, where interactions are directly anchored to objects within the user’s environment. This means that users can immediately access digital functionalities by selecting an object, without the need to navigate through an app or input additional information. By leveraging advanced computer vision and spatial understanding techniques, our AOI framework enables users to seamlessly engage with the physical world as if it were a digital interface.
The object-centric approach offers several advantages over app-centric models. Firstly, it provides a more natural interaction flow, as users can directly engage with objects in their surroundings without the cognitive burden of switching between the physical world and a digital app. Secondly, it minimizes the operational steps required to access digital functionalities, streamlining the user experience and reducing friction. Finally, by anchoring interactions directly to objects, our AOI framework fosters a more immersive and seamless XR experience.
3.1.2. World-Space UI vs. Screen-Space UI
The choice between implementing a world-space UI versus a screen-space UI was informed by our aim to maintain spatial consistency and enhance user engagement with the XR environment. A screen-space UI, fixed relative to the user’s viewpoint, could potentially obfuscate the immersive experience by detaching digital interactions from their physical context. Conversely, our adoption of a world-space UI, where digital elements are anchored to physical objects (akin to ”billboards” in 3D graphics, i.e., user-facing 2D planes in a 3D space), ensures that interactions remain contextually grounded within the user’s real-world environment. We hope to minimize cognitive load by preserving spatial orientation and also leverage the natural human capability to navigate and interact with 3D spaces.
3.1.3. Signaling Identified XR-Objects
To mitigate visual clutter, a common issue in densely populated AR environments, we introduce the use of semi-transparent spheres, or ”bubbles,” as minimalist indicators of interactable objects. This design choice is based on the principle of minimalism and unobtrusiveness, ensuring that users are not overwhelmed by excessive digital information overlaying their physical surroundings. Bubbles serve as subtle prompts that an object is interactive to balance informational availability with spatial aesthetics.
3.1.4. Fixed Number of Top-Level Categories and Actions
The decision to implement a fixed number of top-level categories and actions within the system’s UI was driven by considerations of usability and cognitive efficiency. Limiting the choice set helps mitigate decision fatigue and simplifies the interaction process, making it easier for users to navigate the system’s functionalities. This design philosophy aligns with the Hick-Hyman Law (Wu et al., 2017), which states that increasing the number of choices proportionally increases decision time. By streamlining the number of options available, we are also able to adopt a radial menu with constant reach distance (Samp and Decker, 2010; Pourmemar and Poullis, 2019) instead of dropdown lists, and we facilitate quicker user decision-making and enhance the overall user experience.
In summary, we aim to deliver a seamless, efficient, and immersive XR experience by opting for an object-centric interaction model, employing a world-space UI, utilizing visual bubbles for indicating interactability, and limiting the complexity of user choices.
Figure 5. XR-Objects  instantiates a dedicated MLLM instance for identified object in the scene. Object comparisons are executed by stitching together the relevant objects in the scene before passing the query to an MLLM instance.
[…]
Top: The user’s prompt is supplied to the corresponding object. Each object, represented by a box in this diagram, has an LLM instance. There is also an object comparer running. Bottom: Inside the object comparer, the objects’ images are stitched together to ask the LLM instance a question regarding multiple objects. The correct object is marked green in the real-world scene.
3.2. Categories of Actions
Our system facilitates fluid interactions with a single or multiple objects and enables users to take various digital actions, such as querying real-time information, asking questions, sharing the objects with contacts, or adding spatial notes. Inspired by sub-menus in traditional context menus on desktop computing, we categorized our seven implemented actions into four categories, which we list below.
  1. (1)
    Information: provide an overview; ask a question
  2. (2)
    Compare: ask to compare multiple objects within the view
  3. (3)
    Share: send object to a contact; add to shopping list
  4. (4)
    Anchor: notes; timer; countdown
In the above list, the first two categories (Information and Compare) represent traditional Visual Question Answering (VQA) tasks, while the other two (Share and Anchor) represent traditional widget tasks. We open-source our code on GitHub [^6] XR-Objects open-source project: https://github.com.anonymized and anticipate that the list of integrated actions will be extended in the future by the XR community.
3.3. System Architecture
The implementation of XR-Objects  involves a series of steps to augment real-world objects with functional context menus as illustrated in Figure  4. These steps can be categorized as (1) detecting the objects, (2) localizing and anchoring onto the object, (3) coupling each object with an MLLM for metadata retrieval, and (4) on user input, executing the actions and displaying the output. We use Unity and its AR Foundation [^7] Unity AR Foundation: https://unity.com/unity/features/arfoundation to bring the necessary components for these steps together to build our system. Below, we detail the components and processes that constitute this framework.
3.3.1. Object Detection and Classification
The foundation of XR-Objects  is its robust object detection module, which leverages the capabilities of the Google MediaPipe library (Lugaresi et al., 2019). This module employs a convolutional neural network (CNN) optimized for mobile devices, providing on-device and real-time classification of objects within the user’s camera feed. The system detects objects by providing a class label (e.g., “bottle”, “monitor”, “plant”) and generating 2D bounding boxes, which serve as preliminary spatial anchors for subsequent AR content. The current CNN model is based on the COCO dataset (Lin et al., 2015) which provides 80 labels.
To prioritize user privacy and data efficiency, XR-Objects  further processes only those object regions that are of relevance to the user’s current interaction context. For instance, even though MediaPipe inherently also identifies people in the scene, a region classified as a “person” by the on-device model is not sent to the MLLM-based cloud query system to preserve the privacy of users in the surroundings. Similarly, other classes of objects that are relevant or irrelevant to the user can be customized depending on the AR application (e.g., a plant species search AR app could only run queries for the “plant” class).
3.3.2. 3D Localization and Anchoring
With the object identified, XR-Objects  proceeds to generate AR menus. These menus are spatially anchored to the objects using a combination of the initial 2D bounding boxes and depth information of the scene. We us raycasting to translate the 2D object locations into precise 3D coordinates.
In our system, we used the depth map on the phone (Du et al., 2020) generated through depth from motion (Valentin et al., 2018) by ARCore. Because the location returned by the object detector from the previous step is in 2D screen space, we raycast from this point toward the depth map to “hit” the object and find the corresponding 3D object location in world space, as shown in Figure 4.
At the computed 3D location, we instantiate our object proxy template, which was developed as a prefab (i.e., fully configured GameObjects saved in the AR project for reuse) in Unity. The AR engine ensures that the object proxy stays anchored even when the user changes their view angle.
The object proxy contains the object’s context menu, however, before the user selects the object, it shows up as only a small semi-transparent sphere on the object, which signals to the user that the object has been recognized by the system. Only when the user taps this sphere, the full context menu is shown, otherwise, the menu remains hidden to avoid visual overload.
Our algorithm also includes extra steps to ensure the object proxies do not get spawned at undesired locations or get erroneously duplicated for the same object.
3.3.3. Coupling Each Object with a MLLM
We couple a MLLM with each identified object; thus, we run one LLM conversation instance per object, as shown in Figure 5. We use the cropped bounding box from the first step as the visual input to the MLLM. We also store the conversation history by referring to a conversation ID internally. This object-specific approach enables the MLLM to provide detailed information about the object that extends far beyond the capabilities of traditional object classifiers. For example, it can furnish the object with a wide array of data, including but not limited to product specifications, historical context, and user reviews. As demonstrated in Figure 4, the system is capable of recognizing an object as ”Superior Dark Soy Sauce,” rather than merely identifying it as a ”bottle”—the generic label typically assigned by standard object detection processes in the preceding step.
For our MLLM, we use PaLI (Chen et al., 2023), which runs in the cloud and takes as input the captured region of interest (i.e., object’s cropped image). The MLLM system is able to simultaneously query the Internet (via Google Search) to retrieve additional metadata about the object, e.g., prices and user ratings in the case of a product.
3.3.4. Menu Interaction
Interactions within XR-Objects  are facilitated through a multimodal interface, supporting both touch and voice inputs. This flexible interaction model allows users to engage with the system in a manner that best suits their preferences and the current context. For voice interactions, the system incorporates a speech recognition engine [^8] Speech Recognizer: https://github.com/EricBatlle/UnityAndroidSpeechRecognizer, which enables the processing of natural language commands and queries. As the feedback mechanism, certain user actions, such as selecting a menu option or asking a question, are reflected in the panel overlaid on the object.
When the object is selected by the user, the actions, described in the previous section, are shown. Once an action’s button is tapped, the interaction starts in a panel overlaid on the object.
Information retrieval
For the actions that retrieve real-time data (e.g., getting the answer to the user’s question), we use the object’s MLLM instance. For instance, when the ”info” button is selected, the MLLM-returned object summary is shown. We use the following prompt to create an object summary:
Provide the information from the following list that makes sense for this object. Fill in the missing “…” using info from the Internet. Exclude the one that are irrelevant. Divide the relevant ones with a “*”. * Price: … (give price+vendor+score/ rating) * Cheaper alternatives: name - price * Main ingredients: … (top 2) * Calories: … * Allergens: … * Instructions: … (short) * Care: …(if fashion/tool/plant). Use extremely short answers and exclude answers that are ‘None’ or ‘n/a’ or ‘irrelevant’. Limit to 30 words.
If the user wants to ask a more specific question, they can tap the “ask a question” button, and directly speak out a question.
Object comparisons
For the object compare functionality, we use a dedicated “object comparer” method, which allows us to compile multiple identified objects’ information and provide the combined result as input to a dedicated MLLM instance. As shown in Figure 5, the object comparer stitches all objects’ images together and provides its MLLM instance when the user asks a question $prompt_{user}$ about multiple objects using the “compare” button. The returned response is shown to the user.
If the user’s prompt is a “which” question, the object comparer also executes a follow-up MLLM query under the hood to help visualize the results for this “filtering” type of user question. For this reprompting, we augment the user’s prompt $prompt_{user}$ with a sub-prompt $prompt_{indexing}$ such as:
Considering that the items are ordered from left to right with the first object being index 0, tell me ONLY the correct indices, written as numbers.
Thus, the MLLM returns only the right indices, which we use to mark the relevant objects in the AR view as shown in the bottom-right screenshot in Figure 5.
4. Evaluation
We conducted a user study comparing XR-Objects to a state-of-the-art MLLM assistant interface (Gemini app [^9] Google Gemini: https://play.google.com/store/apps/details?id=com.google.android.apps.bard), referred to as ”Chatbot” from here on, for contextual information retrieval and object-centric interaction. Participants were asked to perform a number of timed VQA tasks and widget interactions in simulated grocery shopping and at-home scenarios, and provided feedback on their experience with each interface through a survey. This study was approved by Institutional Review Board at [anonymized].
4.1. Participants
We recruited 8 participants (6 male, 1 female, 1 preferred not to disclose) between the ages of 25-45. All were fluent English speakers (4 native), were regular shoppers, and all but 1 had used smartphone-based AR at least once.
4.2. Scenario & Tasks
We designed a scenario consisting of a simulated grocery shopping experience (Figure 6a) followed by an at-home experience (Figure 6b) in which users complete a set of 6 tasks using either XR-Objects  or Chatbot. The task categories included: (T1) 2-object Search, (T2) N-object Search, (T3) Share Object, (T4) Get Info (Single Object), (T5) Set/Anchor Timer, and (T6) Create/Anchor Note. We created two variants of this scenario (A and B), each with the same categories but slightly varied tasks. The full set of tasks is presented in Table 1.
Table 1. Task descriptions given in the user study.
For XR-Objects , participants were instructed to use the Compare feature for T1 and T2, the Share feature for T3, the Info feature for T4 and the Anchor feature for T5 and T6. For Chatbot, participants were instructed to take/upload a photo along with a query (using their preferred method of text or voice) to the chat for T1, T2, and T4. For the remaining tasks, participants were told to use Chatbot as if it were connected to a smartphone assistant.
4.3. Procedure
Participants were first given a brief introduction to the study, provided informed consent, and filled out a demographics survey. The experimenter then walked through the functionality of both XR-Objects  and Chatbot, and participants then completed a set of sample tasks on an object not included in the study.
Figure 6. User study setup with mock grocery store (a) and at-home (b) environments. Examples of using XR-Objects  in each case are shown in (c) and (d), respectively.
[…]
(a) Shelves wıth multiple products. (b) A small table with a mug and a can. (c) A phone shows the objects on the shelf in the UI. (d) The same UI looking at the objects on the table.
Each participant completed the tasks in both scenarios A and B. The ordering of both condition (XR-Objects,  Gemini) and scenario (AB) were counterbalanced between participants to prevent ordering effects. Once the tasks were completed, participants were free to use the tool freely for up to five additional minutes however they chose. Participants then completed a post-condition survey on their qualitative experience (Appendix A.1).
After a two minute break, this was repeated with the remaining condition and scenario. Upon completing the tasks with both conditions, participants completed a final survey comparing interactions with XR-Objects  and Chatbot (Appendix 4.4.3). Overall, the study took approximately 45 minutes.
4.4. Measures
4.4.1. Task Completion Time
We recorded the time required to complete tasks T1-T6 for each condition as a measure of overall system performance.
4.4.2. Post-Condition HALIE Survey
Following the completion of all tasks in a given condition (XR-Objects, Chatbot), participants completed a survey evaluating their interactions, adapted from the Human-AI Language-based Interaction Evaluation (HALIE) framework proposed by Lee et al. (Lee et al., 2022). Participants rated their agreement with the following statements (among others) on a 5-point Likert scale. Due to space constraints, the complete survey is provided in Appendix A.1.
4.4.3. Post-Study Form Factor Survey
While the user study was conducted on a smartphone due to limitations of head-mounted display (HMD) camera access, our vision for XR-Objects  is for it to run entirely on the HMD. Therefore, we conducted a post-study survey in which participants envisioned interacting with XR-Objects  on an HMD (e.g., Apple Vision Pro). The survey questions (provided in Appendix A.2) were based on HALIE, but formulated as a forced-choice comparison between the Chatbot and XR-Objects  interaction paradigms.
4.5. Analysis
4.5.1. Task Completion Time
We analyze completion time using traditional t-tests, and confirm normality via Shapiro-Wilk Test.
4.5.2. Distribution Skewness
Skewness ($\gamma_{1}$) quantifies the asymmetry of a given distribution’s shape. For a normal distribution, values of $|\gamma_{1}|$ 0.5 indicate an approximately symmetric distribution. Values of 0.5 ¡ $|\gamma_{1}|$ 1 suggest moderate skewness, while $|\gamma_{1}|$ 1 suggests a highly skewed distribution. This statistical approach, in contrast to visual methods like histograms, is particularly useful for analyzing distributions in Likert-scale questionnaires (King et al., 2018).
4.5.3. Generalized Linear Model
To analyze the data derived from our forced-choice questionnaires, we use a Generalized Linear Model (GLM) based on the Logit Binomial distribution. Unlike regular linear models, GLMs enable regression beyond Gaussian distributions. Considering this data follows a Bernoulli distribution (i.e., datapoints are 0 or 1), our GLM is effectively a log-odds model.
4.6. Results
4.6.1. Time
On average, participants using XR-Objects  required significantly less time (M=217.5 s, SD=58 s) to complete all tasks when compared to the participants using Chatbot (M=286.3, SD=71 seconds), illustrated in Figure 7 and confirmed by a paired t-test (t=-2.8, df=5, p=0.01). This translates to a roughly 31% in task completion time on average for Chatbot users compared to XR-Objects.
Figure 7. Comparison of task completion times between XR-Objects  and Chatbot. The box plot visualizes the distribution of data, individual data points, within-subject comparisons (n=6), and the distribution for each condition.
[…]
The boxplot shows that Chatbot took longer than XR-Objects. The individual datapoints, as well as population density half-violin plots are also shown for comparison.
4.6.2. HALIE Survey
We analyze the HALIE survey results using both traditional non-parametric tests for ordinal data and skewness calculations to assess the distribution of responses (Figure 8).
Figure 8. Likert-scale results of the HALIE survey.
[…]
The HALIE survey rated helpfulness, enjoyment, satisfaction, and responsiveness. For ease, 4 conditions were rated: retrieval, 2 comparison, N comparison, and tool.
While we find no significant differences between Chatbot and XR-Objects on any factor of the HAILE survey (Wilcoxon Paired tests), we find that both approaches of MLLM-enabled real-world search (either via Chatbot or XR-Objects) appear positively rated. Thus, we proceed with a skewness analysis. The most significant skewness of the questionnaire data were found on the questions regarding Ease of Use, and Satisfaction. In particular, Ease of Information Retrieval showed both conditions were highly skewed: XR-Objects  ($\gamma_{1}=1.19$) Chatbot ( $\gamma_{1}=1.8$), making a strong case for MLLM-enabled information retrieval in any form.
Further exploration shows that the responses for Tool Ease were highly skewed for Chatbot ($\gamma_{1}=2.25$). However, those same skews weren’t sustained on the XR-Objects  condition ($\gamma_{1}=0.03$). We hypothesize that this is because XR-Objects  is a research prototype, while the Chatbot used was a fully released product. Nevertheless, we found a moderate skewness on the questionnaire results for the Satisfaction metric only for our prototype XR-Objects  ($\gamma_{1}=0.7$), but not for the Chatbot condition ($\gamma_{1}=0.4$).
4.6.3. Form Factor (Phone or HMD)
The results of the form factor survey are summarized in Figure 9. We applied a GLM with two factors: FormFactor (phone, HMD) and Question (across 12 questionnaire levels), assuming a binomial distribution. Given that the dependent variable’s responses were binary (XR-Objects  or Chatbot), traditional linear models were inadequate as they are tailored to fit Gaussian distributions only. The GLM approach allowed for fitting a Bernoulli distribution and conducting appropriate tests. Our analysis revealed a significant FormFactor effect $F(191,179)=1.917,p<7.05e-08,\eta^{2}=3.8$.
Figure 9. Marimekko-chart mosaic with the form factor survey results.
[…]
The mosaic chart shows a preference for XR-Objects over Chatbot on the HMD condition.
To further assess the model’s effectiveness in predicting Form Factor, we examined the model deviance $(-2LL=209)$ and compared it against the null model’s deviance, which assumed Form Factor was not a consideration $(-2LL=253)$. This comparison demonstrated that our model was more adept at accounting for the variance, where a higher deviance signifies a poorer fit. A chi-square test contrasting the two models yielded a significant difference, with $\chi^{2}=1.6\times 10^{-5},df=12$.
These findings show a clear preference for XR-Objects  in the context of the HMD form factor. Conversely, when using a phone, participants’ preferences between the AI tools (Chatbot or XR-Objects) were split, validating our hypothesis regarding the optimal form factor for tools like XR-Objects.
5. Applications
Through AOI, we envision XR-Objects  to be useful across a variety of real-world applications. By enabling in situ digital interactions with non-instrumented analog objects, we can expand their utility (e.g., enabling a pot to double as a cooking timer), better synthesize their relevant information (e.g., comparing nutritional value), and overall enable richer interactivity and flexibility in everyday interactions. Here we present five example application scenarios from a broad application space we envision (see Figure 10) that highlight the value of XR-Objects.
Figure 10. Envisioned application space for XR-Objects.
[…]
Context: home, work, shopping, nature, sports, mobility. Object relationships: individual, between objects, between object & digital. Persistence: temporary vs. persistent.
5.1. Discovery
XR-Objects  enables users to discover new information about their surroundings by simply pointing their device at objects of interest. As shown in Figure 11, a user points their device at a vase containing different flowers and instantly receive information about each flower type, including its name, price, and care instructions. This on-demand, spatial discovery transforms everyday objects into avenues for understanding or appreciation of everyday objects that normally go unnoticed.
Figure 11. Example of discovery through MLLM interactions on top of XR-Objects.
[…]
A bundle of flowers in a vase. The user can look up different species such as tulip vs. alstroemeria.
5.2. Productivity
For productivity, XR-Objects  could enhance physical documents with digital capabilities such as information retrieval and content anchoring. In Figure 12, a user reading a text book asks how it can be used to solve a particular type of equation, and anchors the response to the textbook for future reference. With added capabilities such as real-time optical character recognition (OCR) to digitize text, users could store and share digital copies of their physical documents for versioning and collaboration.
Figure 12. Interaction mock up of using XR-Objects  for productivity on an HMD with direct touch interactions enabled.
[…]
Two books on a table. The book on the left is about OpenGL. The user points at a page to ask a question.
5.3. Learning
XR-Objects  offers immersive and interactive learning experiences by augmenting physical objects with contextual educational content. By pointing their device at an object, users can access relevant explanations, demonstrations, and quizzes that enhance their understanding of the subject matter.
As illustrated in Figure 13, XR-Objects  can facilitate learning about healthy eating habits for children. A child can point their device at a fruit bowl and instantly see information about the different fruits, such as their names, nutritional values, and the vitamins they contain. This interactive learning experience can help children develop a deeper understanding of the importance of a balanced diet and the benefits of various fruits.
Figure 13. Example of anchoring contextual educational content over real-world objects using XR-Objects.
[…]
The user anchors a spatial note on the fruits: ”a lot of vitamins”.
5.4. IoT Connectivity
XR-Objects’s XR interface is complemented by a MLLM backend, unlike traditional IoT control interfaces that often limit interactions with devices to discrete apps. Through its object tracking, XR-Objects  enables users to interact with their IoT devices in a spatial context, allowing for real-time visual feedback and control within their immediate environment.
For example, in Figure 14, the user controls their smart speaker using XR-Objects. They adjust the volume of the speaker using the touch UI or language commands utilizing the MLLM backend. Such connectivity scenarios expand beyond speakers and could also be enabled on thermostats, smart lights, and other edge IoT devices.
Figure 14. XR-Objects  detects devices and provides custom UI elements for IoT control.
[…]
The user controls the speaker using the AR interface.
5.5. Cooking
Unlike traditional cooking aids, which rely on static recipes or digital screens detached from the cooking environment, XR-Objects  integrates digital intelligence directly into the kitchen, making the cooking process informative and engaging (Figure 1). In our augmented cooking app, as the user places ingredients on the kitchen counter, our system recognizes each item and projects relevant information, such as freshness indicators, nutritional facts, and potential allergens, directly on the ingredients. Users can interact through voice commands or touch elements to ask about potential recipes or to compare ingredients. Using stopwatch or countdown timers, the system embeds the guidance into the cooking space itself. The user can share the final product with their contacts.
6. Discussion
Our daily environment is ubiquitously augmented with various forms of annotations, from product packaging and price tags to traffic signs and personalized post-it reminders. This ubiquitous augmentation, a rudimentary form of augmented reality, is facilitated by the advent of language, writing, and scalable printing technologies. It serves as a means to asynchronously convey contextual information, albeit in a static and limited manner, requiring manual interpretation and application by the user. Although machine-readable markers like barcodes, QR codes, and NFC tags have simplified certain interactions, they fall short in offering dynamic, object-relevant actionable insights.
Augmenting Objects with Intelligence
Advancements in computer vision and LLMs now enable devices to not only recognize generic object categories but also distinguish between individual instances based on their spatial context. This unlocks the potential for personalized, instance-specific interactions to transform objects into intelligent entities with their own ”memories” of past interactions. Accordingly, AOI has the potential to transition from the concept of smart tools to a reality where intelligence is an inherent characteristic of every object.
Context-Aware Interactions
Looking ahead, XR-Objects  could evolve to become even more attuned to the user’s context (e.g. current location and activity, object relationships, persistence of object and data, see Figure 10) to further customize actions and information based on historical interactions with objects in specific settings. For instance, viewing a food item’s packaging in a store might trigger suggestions for understanding its allergens and nutritional content, while the same item at home could offer cooking instructions and allow the user to directly set a digital cooking timer to the appropriate duration. On certain smartphones, for instance, Siri Suggestions [^10] Siri Suggestions on iPhone: https://support.apple.com/guide/iphone/about-siri-suggestions-iph6f94af287/ios already offer context-aware recommendations. The functionality of this sort of context-awareness could be extended using AR interfaces.
Leveraging Emerging Artificial General Intelligence (AGI)
The integration of emerging AGIs (Morris et al., 2024), exemplified by models like Gemini or GPT-4, opens up new opportunities for autonomous problem-solving within XR environments. AGI’s potential to dynamically generate user interfaces and visualizations in response to user queries could transform the way we interact with our physical world. For instance, the user could ask the system to visualize the nutritional values of a product in a pie chart, i.e., by generating the code to create the user interface and graphs on-the-fly without being pre-programmed to create the chart. Going a step further, one may imagine AGI-driven systems that not only respond to user prompts but also proactively offer assistance surfaced through a new action in the context menu, such as assembling a set of Lego blocks into a desired structure through real-time, augmented instructions.
Linking Realities
As we adopt these new innovations in AI and XR, we are likely to see significant changes in how we interact with the physical objects around us. We envision a future where physical items no longer need conventional labels or tags, relying instead on AOI for context and interaction. The merging of digital and physical worlds might bring about new connections, like direct links between digital files and physical items. This could start a new phase where the digital and physical worlds blend together smoothly, without clear boundaries.
7. Conclusion
In this paper, we introduced Augmented Object Intelligence  (AOI), a novel paradigm that seamlessly integrates digital capabilities into physical objects through the use of XR-Objects. Our prototype system demonstrates the potential of AOI to transform how users interact with their surroundings by leveraging advancements in computer vision, spatial understanding, and Multimodal -LLM. The results of our user study show that XR-Objects  significantly outperforms traditional multimodal AI interfaces, with participants completing tasks an average of 24% shorter time and reporting higher levels of satisfaction, ease of use, and perceived responsiveness. By enabling familiar, interactions with everyday objects through anchored AR content and natural language processing, XR-Objects  paves the way for a future where the boundaries between the physical and digital worlds become increasingly blurred. As we continue to expand the capabilities of XR-Objects, we envision a wide range of applications spanning domains such as cooking, productivity, and IoT connectivity, ultimately leading to a more engaging, efficient, and immersive way of interacting with our surroundings.
Appendix A Study Materials
A.1. HALIE Survey
Following the completion of all tasks in a given condition (XR-Objects , Chatbot), participants rated their agreement with the following statements on a 5-point Likert scale, and provided open ended answers to the questions denoted with (*O):
  • H
    (Helpfulness) Independent of its fluency, the AI Tool was helpful for completing my task.
  • *O1
    (Helpfulness) What kinds of aspects did you find helpful or not helpful and why? (Give a concrete example if possible.)
  • J
    (Enjoyment) It was enjoyable using the AI tool to accomplish the tasks.
  • S
    (Satisfaction) Independent of its fluency, I am satisfied with how the AI tool provided its answers.
  • R
    (Responsiveness) Independent of its fluency, I found the AI tool to be a responsive system.
  • E
    (Ease) Overall, it was easy to interact with the AI Tool and accomplish the tasks.
  • ER
    (Ease) Getting information about an object was easy using the AI tool.
  • E2
    (Ease) Comparing two objects was easy using the AI tool.
  • EN
    (Ease) Comparing more than two objects was easy using the AI tool.
  • *O2
    (Change) Did you change how you chose to interact with the AI Tool over the course of the task? If so, how?
  • *O3
    (Description) What adjectives would you use to describe the AI Tool?
  • *O4
    (Compare) How did this interaction compare to your regular in-person shopping and at-home task experiences?
A.2. Form Factor Survey
For each question, participants selected either Chatbot or XR-Objects. They did this questionnaire twice, once we were asking participants to think XR-Objectswas going to run on a Phone, the second time thinking XR-Objectswould run on a headset form factor. This questionnaire included the adapted HALIE questions (H, J, S, R, ER, E2, EN) as well as an additional set of questions detailed below:
  • EC
    (Ease Communications) Sending a message about one of the grocery items would be easier on headset/phone using:
  • ET
    (Ease Timer) Setting a timer would be easier on headset/phone using:
  • EN
    (Ease Note) Creating a note (e.g., reminder to buy more juice) would be easier on a headset/phone using:
  • IS
    (Improved Shopping) Which AI Tool running on a headset/phone would represent a better change compared to your current experience with (in-person) shopping?
  • P
    (Preference) Overall, which AI tool would you prefer on headset/phone?